OpenStudy (anonymous):

Solve the equation for z and show your work. 8(z – 3) = -24

OpenStudy (anonymous):

@pjustine234 8(z - 3) = -24 8z - 24 = -24 8z - 24 + 24 = -24 + 24 8z = 0 8z/8 = 0/8 z = 0 <- answer

OpenStudy (anonymous):

You're amazing. Thanks :) Mind helping me with another problem that's similar?

OpenStudy (anonymous):

welcome! and sure

OpenStudy (anonymous):

Solve the equation for "a" and show your work. 13a – (-76) = -119

OpenStudy (anonymous):

13a - (-76) = -119 13a + 76 = -119 13a + 76 - 76 = -119 - 76 13a = -195 13a/13 = -195/13 a = -15 <- answer

OpenStudy (anonymous):

Thank you!!!:)

OpenStudy (anonymous):

welcome <3